Circular No. 8099

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003cg IN NGC 3169 R. Kotak and W. P. S. Meikle, Imperial College, London; and F. Patat,

European Southern Observatory, on behalf of the European Supernova Collaboration, report that a spectrum (range 427–686 nm) of SN 2003cg (cf. IAUC 8097), obtained at the William Herschel Telescope (+ ISIS) by G. Nelemans on Mar. 22.9 UT, shows it to be a highly-reddened ‘normal’ type-Ia supernova, a few days before maximum (compared to SN 1994D; Patat et al. 1996, MNRAS 278, 111), with a total extinction of A_{v} ∼ 2.5 magnitudes. Very strong, multicomponent Na I D interstellar absorption lines are present in the rest frame of the host galaxy.

Further to IAUC 8091, T. Matheson et al. write that a spectrum of SN 2003cg, obtained by L. Macri on Mar. 25.23 UT, shows it to be a highly reddened, possibly subluminous type-Ia supernova near maximum. Adopting the NED recession velocity (1238 km/s) for the host galaxy, the supernova expansion velocity is 11500 km/s for Si II (rest 635.5 nm). Several titanium absorption features appear stronger than usual, with the entire spectrum resembling that of the subluminous SN 1999by just before maximum (see Garnavich et al. 2003, Ap.J., in press; http://xxx.lanl.gov/abs/astroph/?0105490). The spectrum shows a strong narrow Na I D absorption at the velocity of the host galaxy with an equivalent width of 0.55 nm. There are several other narrow absorptions in the spectrum that appear to coincide with known diffuse interstellar bands (e.g., 544.9, 578.0, 628.3 nm, etc.) at the host velocity. In addition, the overall shape of the spectrum declines toward the blue, implying that the supernova is heavily extinguished by dust. The Schlegel et al. (1998, Ap.J. 500, 525) dust maps indicate a low value for Galactic reddening of E(B − V) = 0.031.

Corrigenda. On IAUC 8097, lines 2, 7, and 16, for Itagati read Itagaki

COMETS C/2002 V6 AND C/2002 V7 (SOHO) Further to IAUC 8073, D. Hammer reports measurements for two addi-

tional Kreutz sungrazing comets found on SOHO C2 website images by K. Černis (C/2002 V6) and R. Matson (C/2002 V7); C/2002 V6 was also visible with the C3 coronagraph. The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2003-F14.

Comet  2002 UT  α_{2000}  δ_{2000}
C/2002 V6 Nov. 13.179  15^{h}08.^{m}2  −21^{o}00′
C/2002 V7  15.850  15 21.5  −20 07
2003 March 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green